OpenStudy (anonymous):

Simplify (3x^2-2x+1)-(x^2-2x-3)+(4x^2-x+2) A.) 6x^2-x+6 B.) 6x^2-5x+6 C.) 6x^2-x D.) 8x^2-5x+6

OpenStudy (anonymous):

All you do is add up all the coefficients with equal powered x's

OpenStudy (anonymous):

u have to combine like terms first

OpenStudy (anonymous):

write it like this: 3x^2-2x+1-x^2+2x+3+4x^2-x+2 collect like terms.

OpenStudy (anonymous):

3x^2-2x+1-x^2+2x+3+4x^2-x+2 =6x^2-x+6

OpenStudy (anonymous):

Not suppose to give out answers.

OpenStudy (anonymous):

idk how to explain , sorry

OpenStudy (anonymous):

Would you do 3x^2-x^2 and get 2x^2, and then add the 4x^2 to get 6x^2?

OpenStudy (anonymous):

yes but its more safe if u add the all negative and positive before you add and minus like 4x^2+ 3x^2 = 7 x^2 before u minus the x^2= and u can get 6x^2

OpenStudy (anonymous):

Ohh okay, I get it now. Thanks ! :)

OpenStudy (anonymous):

:)